Suppose that u and v are unit vectors such that (u, v) =-1/2. Find |3u+v|.
Answer #1

|3u+v|^2=< 3u+v,3u+v>=<3u,3u+v>+<v,3u+v>

=<3u,3u>+<3u,v>+<v,3u>+<v,v> =9|u|^2+3<u,v>+3<v,u>+|v|^2=<3u,3u>+<3u,v>+<v,3u>+<v,v> =9|u|^2+3<u,v>+3<v,u>+|v|^2 =9+6<u,v>+1 ==9+6<u,v>+1 =9+6*(-1/2)+1 =9-3+1=7

So |3u+v|=\sqrt 7

Above we are using properties of inner product.